# DAMS-DETR: Dynamic Adaptive Multispectral Detection Transformer with Competitive Query Selection and Adaptive Feature Fusion

### 04 Mar 2024 v2

@@ Abstract
Infrared-visible object detection aims to achieve robust even full-day object detection by fusing the complementary information of infrared and visible images. However, highly dynamically variable complementary characteristics and commonly existing modality misalignment make the fusion of complementary information difficult. In this paper, we propose a Dynamic Adaptive Multispectral Detection Transformer (DAMS-DETR) based on DETR to simultaneously address these two challenges. Specifically, we propose a Modality Competitive Query Selection strategy to provide useful prior information. This strategy can dynamically select basic salient modality feature representation for each object. To effectively mine the complementary information and adapt to misalignment situations, we propose a Multispectral Deformable Cross-attention module to adaptively sample and aggregate multi-semantic level features of infrared and visible images for each object. In addition, we further adopt the cascade structure of DETR to better mine complementary information. Experiments on four public datasets of different scenes demonstrate significant improvements compared to other state-of-the-art methods. The code will be released at https://github.com/gjj45/DAMS-DETR.
@@ 1 Introduction
However, these methods tend to neglect the modality interference encountered in complex scenes during the fusion process. For the case that the object signal in one modality is poor or absent, directly fusing the information of two modalities will bring in unuseful interference information, which could lead to feature confusion and thus degrade the object detection performance. For example, as shown in Fig. 1(a), the pedestrian within the smoke fully disappears, and intuitively, the best way is to suppress or discard the visible information for that pedestrian. Some works learn a global fusion weight to adapt to specific scenes, the representative ones of which are to adopt the illumination-aware network to obtain illumination score as the global fusion weight [12, 41, 33]. Other works learn local region fusion weights through bounding-box level semantic segmentation [12, 41, 33], or regions of interest (ROI) prediction [40, 39, 9].
@@
In this paper, we propose a novel adaptive infrared-visible object detection method based on DETR [42, 35, 19]. Inspired by the deformable cross attention [42], we propose a Multispectral Transformer Decoder with Multispectral Deformable Cross-attention module to simultaneously address the above two challenges. Specifically, we adopt an effective strategy of adaptive sparse feature sampling and weight aggregation on two-modality feature maps of different semantic levels. This strategy can effectively fuse fine-grained complementary information even when two modalities are misaligned. Since the two challenges of fine-grained information fusion and modality alignment are simultaneously handled in a single module, our method is more efficient than existing methods which usually handle them separately. Furthermore, unlike the one-step fusion strategy adopted by existing methods, the information fusion of each specific object in our method happens on different semantic levels, which makes the complementary information fully be mined and utilized. Actually, we observed that the complementary information of two modalities also dynamically varies with the semantic levels, as discussed in Sec. 3.3. This is similar to our observation on scenes and objects discussed previously. Thus, our adaptive multi-level fusion is more reasonable.
In order to provide reliable input at the early stage, we design a Competitive Query selection strategy to select dominant modality features for each object as initial position and content queries to build a basic salient feature representation for the Multispectral Transformer Decoder, which can provide useful prior information for following processing. To further exploit more reliable and comprehensive complementary information step by step, the cascaded layer structure of DETR is employed in this paper. Totally, our method is similar to the human observation pattern which dynamically focuses on objects in each modality and gradually aggregates key information of two modalities.
@@
  • 1.
    We propose a novel infrared-visible object detection method, named DAMS-DETR, which can dynamically focus on dominant modality objects and adaptively fuse complementary information.
  • 2.
@@ 3.1 Overview
The overview of our DAMS-DETR is illustrated in Fig. 2. Our method contains four main components: two modality-specific CNN backbones, two modality-specific Efficient Hybrid Encoders, Modality Competitive Query Selection, and a Multispectral Transformer Decoder.
Figure 2: Overview of DAMS-DETR. Our DAMS-DETR comprises four main components: two modality-specific CNN backbones to extract features, two modality-specific Efficient Hybrid Encoders [19] to encode features, a Modality Competitive Query Selection module to select initial object queries, and a Multispectral Transformer Decoder to mine complementary information.
Given a pair of infrared and visible images, we first extract and encode their features separately using two modality-specific CNN backbones and two modality-specific Efficient Hybrid Encoders. Subsequently, the encoded features are flattened, concatenated, and input into the Modality Competitive Query Selection module. This module selects salient modality features to serve as the initial object queries. Next, these modality-specific object queries enter the Multispectral Transformer decoder, which refines them with multiple semantic levels of infrared and visible feature maps through cascaded decoder layers. Finally, these refined object queries are mapped through the detection head to obtain the bounding boxes and classification scores of all objects.
The Efficient Hybrid Encoder we adopted is from RT-DETR [19]. This encoder combines Transformer and CNN to significantly reduce computational complexity. In the following, we will elaborate on the proposed Modality Competitive Query Selection strategy and Multispectral Transformer Decoder with the Multispectral deformable cross-attention module in detail.
@@ 3.3 Multispectral Transformer Decoder
where $d\in\{2,3,\ldots,D\}$, $MLP$ consists of two linear projection layers, $\sigma$ represents sigmoid function, $\sigma^{-1}$ represents inverse sigmoid function, and $b_{q}^{1}$ is the initialized anchor boxes. The initial anchor setting is consistent with the two-stage method in Deformable DETR [42]. This refined 4D reference point serves as the reference position constraint for subsequent sampling on multi-semantic infrared and visible feature maps in the Multispectral Deformable Cross-attention module.
@@
Figure 5: Visualization of feature sampling at different semantic levels in different decoder layers. We use different colors of points to represent the results of sampling in different semantic layers, where blue, green, and red represent sampling points on low-level, middle-level, and high-level semantic features maps respectively. Brightly colored and large points indicate relatively high attention weights. (a) The green boxes in the visible image represent aligned bounding boxes, which show the sampling points in each modality are concentrated on the right instance locations. (b) Objects occluded by background and smoke tend to predominantly focus on the infrared modality at subsequent decoder layers. (c) Objects in good illumination conditions and those less distinguishable in the infrared modality tend to predominantly focus on the visible modality at subsequent decoder layers.
@@
where $m\in\{1,2\}$ denotes the visible and infrared modalities, $h$ indexes the attention head, $l$ indexes the input feature semantic level, and $k$ indexes the sampling point. $A_{mhlqk}$ and $\Delta\mathrm{p}_{mhlqk}$ denote the $k$-th attention weight and sampling point in the $l$-th feature semantic level and the $h$-th attention head within the $m$ modality, respectively. The attention weight $A_{mhlqk}$ is normalized by $\sum_{m}\sum_{l=1}^{L}\sum_{k=1}^{K}A_{mhlqk}=1$. Function $\Phi_{l}\left(\hat{p}_{q}\right)$ scales $\hat{p}_{q}$ to the $l$-th semantic level feature map and function $\Psi_{l}\left(\Delta\mathrm{p}_{mhlqk}\right)$ constrains the predicted offset within the range of $b_{q}$, so as to focus on the information around the object and reduce the difficulty of optimization.
@@ 4.2 Implementation Details
We employ ResNet50 [6] as the backbone for both the infrared and visible branches, with a feature map semantic level of L = 3. The Efficient Hybrid Encoder contains one layer, while the MS-Decoder contains six layers. We set the number of attention heads, sampling points, and selected queries as H = 8, K = 4, and N = 300, respectively.
@@ 4.3 Comparison with SOTA
Comparisons on VEDAI: As shown in Tab. 4, our method still can work well for small objects, and outperforms most of the SOTA methods. Especially, on $m$AP50, our method outperforms the best (LRAF-Net [5]) by 5.6%. On $m$AP, our method does not outperform LRAF-Net [5] which is a CNN-based detector. The reason could be that in comparison to long-range information extraction, small object detection more relies on local feature extraction, and CNNs could be more advantageous than Transformer in this regard. On the other hand, the smaller bounding boxes of these small objects are sensitive to displacements, and even a few pixel displacements will lead to significant changes in IOU and subsequent $m$AP reduction. Therefore, high precision in bounding box prediction is crucial for small object detection. Relatively, CNN methods can achieve higher bounding box accuracy, resulting in a higher $m$AP.
@@
Nevertheless, the improvement in $m$AP50 demonstrates the potential of our method in small object scenes.
@@
Effect of Content Query Selection (CQS). We analyze the performance of different content query strategies in infrared-visible object detection. For comparison, we take the single branch of our network as the single-modality detection method. In the 8th and 9th rows of the Tab. 5, without CQS means setting the content query as a learnable query, while the position query is still obtained by Modality Competitive Query Selection (similar to mixed query selection in DINO [35]). We can see that using CQS leads to better results, as it provides stronger prior information for subsequent multispectral cross-attention in the Multispectral Decoder. However, as shown in the first 4 rows of Tab. 5, the selection or learnable of content query has a limited impact on single-modality object detection, since the feature representation of single modality is relatively consistent.
@@ 4.5 Limitation
Our method is built upon DETR, so the accurate detection performance on small objects may not be as competitive as other CNN-based detectors, as shown in Tab. 4. This is attributed to the fact that the transformer model prioritizes global information, but it is less effective in extracting local information compared to the CNN model, resulting in less precise boundaries for small object bounding boxes. Additionally, our method is effective in addressing common misalignment in the majority of cases. However, it may not handle extreme misalignment situations well, as the object features of the modality will be lost when objects exceed the range of the 4D reference point.
@@ 5 Conclusion
In this paper, we propose DAMS-DETR to simultaneously address complementary information fusion and modality misalignment problems in infrared-visible object detection. Through Modality Competitive Query Selection, DAMS-DETR can dynamically choose salient modality feature representations for specific objects based on complementary characteristics. In the Multispectral Deformable Cross-attention module, we link feature fusion and modality misalignment to mine reliable complementary information at multi-semantic levels. Experiments on four datasets with different scenarios demonstrate that the proposed method achieves significant improvements compared to other state-of-the-art methods. We hope this work can inspire more DETR-based infrared-visible object detection work.

### 07 Mar 2024 v3

# DAMSDet: Dynamic Adaptive Multispectral Detection Transformer with Competitive Query Selection and Adaptive Feature Fusion

@@ Abstract
Infrared-visible object detection aims to achieve robust even full-day object detection by fusing the complementary information of infrared and visible images. However, highly dynamically variable complementary characteristics and commonly existing modality misalignment make the fusion of complementary information difficult. In this paper, we propose a Dynamic Adaptive Multispectral Detection Transformer (DAMSDet) to simultaneously address these two challenges. Specifically, we propose a Modality Competitive Query Selection strategy to provide useful prior information. This strategy can dynamically select basic salient modality feature representation for each object. To effectively mine the complementary information and adapt to misalignment situations, we propose a Multispectral Deformable Cross-attention module to adaptively sample and aggregate multi-semantic level features of infrared and visible images for each object. In addition, we further adopt the cascade structure of DETR to better mine complementary information. Experiments on four public datasets of different scenes demonstrate significant improvements compared to other state-of-the-art methods. The code will be released at https://github.com/gjj45/DAMSDet.
@@ 1 Introduction
However, existing methods tend to neglect the modality interference encountered in complex scenes during the fusion process. For the case that the object signal in one modality is poor or absent, directly fusing the information of two modalities will bring in unuseful interference information, which could lead to feature confusion and thus degrade the object detection performance. For example, as shown in Fig. 1(a), the pedestrian within the smoke fully disappears, and intuitively, the best way is to suppress or discard the visible information for that pedestrian. Some works learn a global fusion weight to adapt to specific scenes, the representative ones of which are to adopt the illumination-aware network to obtain illumination score as the global fusion weight [12, 41, 33]. Other works learn local region fusion weights through bounding-box level semantic segmentation [12, 41, 33], or regions of interest (ROI) prediction [40, 39, 9].
@@
In this paper, we propose a novel adaptive infrared-visible object detection method, which contains a Multispectral Transformer Decoder with Multispectral Deformable Cross-attention module to simultaneously address the above two challenges inspired by the deformable cross attention [42]. Specifically, we adopt an effective strategy of adaptive sparse feature sampling and weight aggregation on two-modality feature maps of different semantic levels. This strategy can effectively fuse fine-grained complementary information even when two modalities are misaligned. Since the two challenges of fine-grained information fusion and modality alignment are simultaneously handled in a single module, our method is more efficient than existing methods which usually handle them separately. Furthermore, unlike the one-step fusion strategy adopted by existing methods, the information fusion of each specific object in our method happens on different semantic levels, which makes the complementary information fully be mined and utilized. Actually, we observed that the complementary information of two modalities also dynamically varies with the semantic levels, as discussed in Sec. 3.3. This is similar to our observation on scenes and objects discussed previously. Thus, our adaptive multi-level fusion is more reasonable.
In order to provide reliable input at the early stage, we design a Competitive Query selection strategy to select dominant modality features for each object as initial position and content queries to build a basic salient feature representation for the Multispectral Transformer Decoder, which can provide useful prior information for following processing. To further exploit more reliable and comprehensive complementary information step by step, the cascaded layer structure of DETR [35] is employed in this paper. Totally, our method is similar to the human observation pattern which dynamically focuses on objects in each modality and gradually aggregates key information of two modalities.
@@
  • $\bullet$
    We propose a novel infrared-visible object detection method, named DAMSDet, which can dynamically focus on dominant modality objects and adaptively fuse complementary information.
  • $\bullet$
@@ 3.1 Overview
The overview of our DAMSDet is illustrated in Fig. 2. Our method contains four main components: two modality-specific CNN backbones, two modality-specific Efficient Encoders, Modality Competitive Query Selection, and a Multispectral Transformer Decoder.
Figure 2: Overview of DAMSDet. Our DAMSDet comprises four main components: two modality-specific CNN backbones to extract features, two modality-specific Efficient Encoders [19] to encode features, a Modality Competitive Query Selection module to select initial object queries, and a Multispectral Transformer Decoder to mine complementary information and refine queries.
Given a pair of infrared and visible images, we first extract and encode their features separately using two modality-specific CNN backbones and two modality-specific Efficient Encoders. Subsequently, the encoded features are flattened, concatenated, and input into the Modality Competitive Query Selection module. This module selects salient modality features to serve as the initial object queries. Next, these modality-specific object queries enter the Multispectral Transformer decoder, which refines them with multiple semantic levels of infrared and visible feature maps through cascaded decoder layers. Finally, these refined object queries are mapped through the detection head to obtain the bounding boxes and classification scores of all objects.
The Efficient Encoder combines Transformer and CNN to significantly reduce computational complexity, following the structure of RT-DETR [19]. In the following, we will elaborate on the proposed Modality Competitive Query Selection strategy and Multispectral Transformer Decoder with the Multispectral deformable cross-attention module in detail.
@@ 3.3 Multispectral Transformer Decoder
where $d\in\{2,3,\ldots,D\}$, $MLP$ consists of two linear projection layers, $\sigma$ represents the sigmoid function, $\sigma^{-1}$ represents the inverse sigmoid function, and $b_{q}^{1}$ is the initialized anchor boxes. The initial anchor setting is consistent with the two-stage method in Deformable DETR [42]. This refined 4D reference point serves as the reference position constraint for subsequent sampling on multi-semantic infrared and visible feature maps in the Multispectral Deformable Cross-attention module.
@@
Figure 5: Visualization of feature sampling at different semantic levels in different decoder layers. Different colors of points represent the results of sampling in different semantic layers, where blue, green, and red represent sampling points on low-level, middle-level, and high-level semantic features maps respectively. Brightly colored and large points indicate relatively high attention weights. (a) The green boxes in the visible image represent aligned bounding boxes, which show the sampling points in each modality are concentrated on the right instance locations. (b) Objects occluded by background and smoke tend to predominantly focus on the infrared modality at subsequent decoder layers. (c) Objects in good illumination conditions and those less distinguishable in the infrared modality tend to predominantly focus on the visible modality at subsequent decoder layers.
@@
where $m\in\{1,2\}$ denotes the visible and infrared modalities, $h$ indexes the attention head, $l$ indexes the input feature semantic level, and $k$ indexes the sampling point. $A_{mhlqk}$ and $\Delta\mathrm{p}_{mhlqk}$ denote the $k$-th attention weight and sampling point in the $l$-th feature semantic level and the $h$-th attention head within the $m$ modality, respectively. The attention weight $A_{mhlqk}$ is normalized by $\sum_{m}\sum_{l=1}^{L}\sum_{k=1}^{K}A_{mhlqk}=1$. The function $\Phi_{l}\left(\hat{p}_{q}\right)$ scales $\hat{p}_{q}$ to the $l$-th semantic level feature map and function $\Psi_{l}\left(\Delta\mathrm{p}_{mhlqk}\right)$ constrains the predicted offset within the range of $b_{q}$, so as to focus on the information around the object and reduce the difficulty of optimization.
@@ 4.2 Implementation Details
We employ ResNet50 [6] as the backbone for both the infrared and visible branches, with a feature map semantic level of L = 3. The Efficient Encoder contains one layer, while the MS-Decoder contains six layers. We set the number of attention heads, sampling points, and selected queries as H = 8, K = 4, and N = 300, respectively.
@@ 4.3 Comparison with SOTA
Comparisons on VEDAI: As shown in Tab. 4, our method still can work well for small objects, and outperforms most of the SOTA methods. Especially, on $m$AP50, our method outperforms the best (LRAF-Net [5]) by 5.6%. On $m$AP, our method does not outperform LRAF-Net [5] which is a CNN-based detector. The reason could be that in comparison to long-range information extraction, small object detection more relies on local feature extraction, and CNNs could be more advantageous than Transformer in this regard. On the other hand, the smaller bounding boxes of these small objects are sensitive to displacements, and even a few pixel displacements will lead to significant changes in IOU and subsequent $m$AP reduction.
@@
Therefore, high precision in bounding box prediction is crucial for small object detection. Relatively, CNN methods can achieve higher bounding box accuracy, resulting in a higher $m$AP. Nevertheless, the improvement in $m$AP50 demonstrates the potential of our method in small object scenes.
@@
Effect of Content Query Selection (CQS). We analyze the performance of different content query strategies in infrared-visible object detection. For comparison, we take the single branch of our network as the single-modality detection method. In the 8th and 9th rows of the Tab. 5, without CQS means setting the content query as a learnable query, while the position query is still obtained by Modality Competitive Query Selection (similar to mixed query selection in DINO [35]). We can see that using CQS leads to better results, as it provides stronger prior information for subsequent multispectral cross-attention in the Multispectral Decoder. However, as shown in the first 4 rows of Tab. 5, the selection or learnability of content query has a limited impact on single-modality object detection, since the feature representation of single modality is relatively consistent.
@@ 4.5 Limitation
Our method is effective in addressing common misalignment in the majority of cases. However, it may not handle extreme misalignment situations well, as the object features of the modality will be lost when objects exceed the range of the 4D reference point.
@@ 5 Conclusion
In this paper, we propose DAMSDet to simultaneously address complementary information fusion and modality misalignment problems in infrared-visible object detection. Through Modality Competitive Query Selection, DAMSDet can dynamically choose salient modality feature representations for specific objects based on complementary characteristics. In the Multispectral Deformable Cross-attention module, we link feature fusion and modality misalignment to mine reliable complementary information at multi-semantic levels. Experiments on four datasets with different scenarios demonstrate that the proposed method achieves significant improvements compared to other state-of-the-art methods.